OpenStudy (anonymous):

Determine the coordinates of the point of intersection of the graphs y = log(x-2) and y= 1 - log(x+1).

hartnn (hartnn):

to get the points of intersection, just solve those 2 equations simultaneously. y = log(x-2) and y= 1 - log(x+1). gives log(x-2) =1 - log(x+1). got this ? can you solve for x there ?

hartnn (hartnn):

how ?

OpenStudy (anonymous):

@hartnn x=-3, 4.

hartnn (hartnn):

did u post x=-3 , 4 ?

hartnn (hartnn):

oh yes, those are correct :)

hartnn (hartnn):

so, when y = log (x-2) and x=4, what is y=.. ?

OpenStudy (anonymous):

@hartnn y = log2.

hartnn (hartnn):

now you should note that log of negative numbers is not defined! so, x CANNOT take the value of -3 so, only value of x is 4 and only point of intersection is (4,log 2) makes sense ?

OpenStudy (anonymous):

@hartnn So the answer is (4, log2)?

hartnn (hartnn):

yes, absolutely :)

OpenStudy (anonymous):

@hartnn But the answer stated here is (0, -1).... O.o

hartnn (hartnn):

but x cannot take the value = 0, then log (0-2) = log (-2) is actually not defined. thats surely not the answer...

hartnn (hartnn):

see i verified it online too http://www.wolframalpha.com/input/?i=log%28x-2%29+%3D1+-+log%28x%2B1%29.&a=*FunClash.log-_*Log10- if you see the graph there, it has x=4, and y=log 2

hartnn (hartnn):

the answer has to be 4, log 2 :) there must be some typing error in your answer sheet....

OpenStudy (anonymous):

@hartnn Oh, okay.... Thank you so much!!

hartnn (hartnn):

welcome ^_^